OpenStudy (anonymous):

3|x − 8| + 15 = 18 a) x = 1, x = 17 b) x = −7, x = 9 c) x = 7 x = 9 d) no solution

OpenStudy (anonymous):

since there's an absolute value notation you will have two equations which means there are two solution for x. \[3[-(x-8)]+15=18\] and \[3(x-8)+15=18\] solve values for "x" in each equation.

OpenStudy (anonymous):

is it a?

OpenStudy (anonymous):

nope...

OpenStudy (tkhunny):

3|x − 8| + 15 = 18 Simplify your life!! Divide by 3. |x − 8| + 5 = 6